OpenStudy (anonymous):

What is the maximum or minimum value of 100x^2-1700x? Please help, online test due tonight

OpenStudy (highschoolmom2010):

here at OS it is requested that you dont post test/exam questions :D

OpenStudy (anonymous):

Well, he said I could use every resource I have lol

OpenStudy (highschoolmom2010):

ah well that is alright :D

OpenStudy (anonymous):

could you help me?

OpenStudy (highschoolmom2010):

nope sorry @mathstudent55

OpenStudy (anonymous):

if you take the derivative then set the equation to = 0 then solve you will have a max or min \[\frac{d}{dx} 100x^2-1700x = 200x-1700\]now we need to set F(x) to zero then solve for x \[200x - 1700=0\]\[200x = 1700\]\[x = \frac{1700}{200}\]

OpenStudy (anonymous):

in our case since we are positive we have a min

OpenStudy (anonymous):

and also if you graph the function you can see the min